OpenStudy (anonymous):

9sqrt{y^9} 9sqrt{y^9} @Mathematics

OpenStudy (anonymous):

\[\sqrt[9]y ^{9}{?}\]

jimthompson5910 (jim_thompson5910):

\[\LARGE \sqrt[9]{y^9} = y\]

OpenStudy (anonymous):

do the 9s just cancel out?

jimthompson5910 (jim_thompson5910):

effectively yes for any integer n that is an odd number \[\LARGE \sqrt[n]{x^n} = x\]

OpenStudy (anonymous):

ok i got it thanks.

jimthompson5910 (jim_thompson5910):

np